OpenStudy (anonymous):

Question - 18: Signals And Systems: Find the \(\color{red}{\text{Laplace Transform}}\) of following functions: \(\color{green}{a) \qquad f(t) = 1}\) \(\color{blue}{b) \qquad f(t) = e^{-a|t|}}\)

OpenStudy (anonymous):

Laplace Transform of a function is given as: \[\large \color{green}{F(s) = \int\limits_{-\infty}^{\infty} f(t) \cdot e^{-st} \cdot dt}\]

OpenStudy (anonymous):

For first part: \[\large F(s) = \int\limits_{-\infty}^{\infty} e^{-st} \cdot dt\]

OpenStudy (anonymous):

This integral does not seem to be Converging.. May be, I have to find this by using Inverse Laplace Transform..

OpenStudy (anonymous):

@dan815 @eliassaab

OpenStudy (zarkon):

the integral goes from 0 to infinity

OpenStudy (anonymous):

For Bilateral Laplace, limit is from \(-\infty\) to \(\infty\) no?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

So?

OpenStudy (zarkon):

the integral will not converge then for the 1st one

OpenStudy (zarkon):

you should also use " bilateral Laplace transform" in the title of your question since the regular Laplace transform is an integral on the positive reals

OpenStudy (anonymous):

I want to know both..

OpenStudy (anonymous):

Actually, the thing is, my teacher has taught Bilateral first and then he gave some questions for homework, and then after it, he taught unilateral laplace. May be he was checking that what result we make out from that homework..

OpenStudy (anonymous):

So, for bilateral laplace, it is coming out to be Not-defined..

OpenStudy (zarkon):

for the first one...that is correct

OpenStudy (anonymous):

Let me try for the second one for bilateral laplace.

OpenStudy (anonymous):

\[\implies \int\limits_{- \infty}^{0} e^{at} e^{-st} \cdot dt + \int\limits_{0}^{\infty} e^{-at}e^{-st} \cdot dt\]

OpenStudy (anonymous):

\[\implies \frac{1}{a-s} + \frac{1}{a+s}\]

OpenStudy (zarkon):

you need to put a bound on the possible values of a,s

OpenStudy (anonymous):

\[\frac{-2a}{s^2 - a^2}\]

OpenStudy (anonymous):

For first one, Re(s) < a.. Right?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

For second one, Re(s) > -a

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

\[\frac{-2a}{s^2 - a^2} , \quad -a < \Re\{s\} < a\]

OpenStudy (zarkon):

looks fine

OpenStudy (anonymous):

One more thing, when we use Unilateral, these conditions come or not?

OpenStudy (anonymous):

For second one for unilateral: \[F(s) = \frac{1}{s+a}\]

OpenStudy (zarkon):

the condition for the second integral will still apply

OpenStudy (anonymous):

Here also we need to define ROC?

OpenStudy (anonymous):

Okay, got it.. Thank you @Zarkon it really helped a lot to understand. :)

OpenStudy (anonymous):

One more thing..

OpenStudy (anonymous):

I have a doubt in Fourier transform, can you clarify that?

OpenStudy (zarkon):

I haven't done Fourier analysis in a long time, but I can look

OpenStudy (anonymous):

Suppose you have two functions, that are totally same but the dc value of two are different..

OpenStudy (anonymous):

and you evaluate their Fourier Transform by using Time-Differentiation Property//

OpenStudy (anonymous):

On differentiating, the dc value of both will become zero, and you will get same transform for both, what is the condition to be used here to have different transforms?

OpenStudy (anonymous):

Like one function is : f(t), and other is f(t)+ (1/2).. When you use differentiation property, for second 1/2 will become 0, and you have both functions as same..

OpenStudy (anonymous):

getting my doubt?

OpenStudy (zarkon):

I don't remember. That is something I would have to refresh myself on. Last time I did stuff like this was around 2001

OpenStudy (anonymous):

I know the same condition for Fourier Series but don't know for Fourier Transform..

OpenStudy (anonymous):

Okay, you have all the time.. :)

OpenStudy (anonymous):

2001?? I was like in 5th grade at that time.. :P

OpenStudy (zarkon):

I guess I'm a little older than you ;)

OpenStudy (anonymous):

you sure you are "little" older than me?? :P

OpenStudy (zarkon):

I'm in my 40's

OpenStudy (anonymous):

That is not little for a person who is in his 20's..

OpenStudy (zarkon):

ic ;)

OpenStudy (zarkon):

you were born around the time I graduated from high school :)

OpenStudy (anonymous):

If you are planning to read for my doubt, then do read Time-integration property too..

OpenStudy (anonymous):

I guess it is something to do with that..

OpenStudy (zarkon):

ok...I'll see if I can find something. though I only have my home library and the internet to work from. I'm done in my office till next year.

OpenStudy (anonymous):

yeah for sure.. 1991...

OpenStudy (zarkon):

yep...that is the year I graduated HS :)

OpenStudy (anonymous):

:)

OpenStudy (zarkon):

I was born in 1972

OpenStudy (anonymous):

I am not going to tell you the date otherwise you will say I was graduated on that date. :P

OpenStudy (zarkon):

I doubt that

OpenStudy (anonymous):

It was 3rd June 1991, Monday..

OpenStudy (zarkon):

I don't remember the day anyway.. it was a long time ago :)

OpenStudy (anonymous):

:)

OpenStudy (zarkon):

I'm sure the graduation was not on a Monday

OpenStudy (anonymous):

It was really nice of you for sharing such information.. :)

OpenStudy (anonymous):

yeah we don't graduate on Monday which is just after Sunday. :P

OpenStudy (zarkon):

you too

OpenStudy (anonymous):

Thanks for all your help.. Good day..!! :)

OpenStudy (zarkon):

later